# How Cohesive Are Community Search Results on Online Social Networks?: An Experimental Evaluation

### v1

@@ 1. Introduction
In summary, this work makes the following novel contributions. (a) It is the first study to assess the cohesiveness of social network communities generated by various community search methods, analyzed through the lens of cohesion theories in social psychology. (b) It introduces a set of psychology-informed cohesiveness measures for evaluation, which are unavailable in existing community search evaluation toolkits. (c) The experimental findings presented in this work are novel and have not been reported in previous studies.
@@ 3. The Community Search Problem
Problem 3.1 (Community Search (CS) Problem).
Given a graph $G=(V,E)$ and a set of query nodes $Q\subseteq V$, the community search problem seeks to find a community that contains $Q$ and satisfies the following properties:
@@
Traditional CS algorithms. Generally, traditional CS algorithms formulate a community by specifying its subgraph properties (Fang et al., 2020). The commonly adopted cohesiveness property is structural-based, including k-core (Sozio and Gionis, 2010; Cui et al., 2014; Barbieri et al., 2015; Yao and Chang, 2021), k-truss (Huang et al., 2014; Akbas and Zhao, 2017; Liu et al., 2020; Zhang et al., 2023), k-clique (Yuan et al., 2017), and k-ECC (Chang et al., 2015; Hu et al., 2016, 2017). As such, these metrics are often referred to as structural cohesiveness metrics (structural metrics for brevity). Additionally, studies incorporate attribute information to formalize cohesiveness by refining existing structural metrics or proposing new attribute cohesiveness metrics alongside the structural ones. Examples of the former approach include attributed $pkd$-truss community (Xie et al., 2021), $(k,\theta)$-community (Miao et al., 2022), $(k,p)$-core (Lu et al., 2022), and $\beta$-temporal core (Lin et al., 2024). Examples of the latter include keyword cohesiveness (Fang et al., 2016), spatial cohesiveness (Fang et al., 2017; Chen et al., 2018), and query cohesiveness (Das et al., 2021).
@@ 4.1. Methodology
To measure this concept, a conceptual model has been proposed that identifies two key perceptions binding members to their group: a member’s perception of the group as a whole (labeled as group integration (GI)) and a member’s personal attraction to the group (labeled as individual attractions to group (ATG)). Each perception is further divided into task and social aspects, resulting in four correlated constructs that influence group cohesion: group integration-task (GI-T), group integration-social (GI-S), individual attractions to group-task (ATG-T), and individual attractions to group-social (ATG-S) (Carron et al., 1985). The Group Environment Questionnaire (GEQ) derived from this model consists of 18 items and has demonstrated its reliability and validity in various group contexts (Carron et al., 1985; Blanchard et al., 2000; Anderson and Warner, 2017). The cohesion score for a group is calculated by averaging individual GEQ responses (Carron et al., 2003).
@@ 4.2. Comparison
Table 2 summarizes whether existing measures (introduced in Section 3) capture the cohesion features outlined in the adapted GEQ (Table 1). Observe that both ATG and GI items incorporate underlying structural information, yet in distinct ways. Moreover, ATG places greater emphasis on personal opinions, as they carry more psychological insights that influence a user’s membership. Items are labeled “S” for structural information and “P” for psychological information. $\checkmark$”, $\triangle$”, and “$\times$” denote that an item is fully captured, partially captured, or not captured by existing measures, respectively.
@@ 5. Psychology-informed Cohesiveness Measures
5.1. Enjoyment Index (EI)
Item 1 emphasizes the user’s enjoyment level during social interactions. Enjoyment is often measured concerning positive emotions or needs fulfillment during interactions (Syn and Oh, 2015; Kawamichi et al., 2016; Lin et al., 2009). Accordingly, we identify the enjoyment index as a cohesiveness indicator, quantified by the cumulative sentiments elicited during interactions.
@@ 5.1. Enjoyment Index (EI)
Let $\mathcal{E}_{H,i\cdot}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},j\neq i,t\leq t_{cur}\}$ denote the set of activities where user $u_{i}$ interacts with other group members up to time $t_{cur}$. The decay function $\Phi(\cdot)$ is also introduced to assign greater weight to recent activities that better capture user interests (Sarker et al., 2019). Thereby, we define the enjoyment index as follows.
@@
5.2. Sentimental Interaction Tendency (SIT)
Now consider Item 5 in Table 1. Having best friends in the group indicates cohesiveness, prompting the need to measure friendships among group members. Friendship generally develops through the exchange of information (Chan and Lo, 2014) and is characterized by reciprocity, closeness, and emotional attachment (Amichai-Hamburger et al., 2013; Paxton and Moody, 2003). To this end, we propose sentimental interaction tendency to characterize friendships developed through cumulative information exchange.
@@ 5.2. Sentimental Interaction Tendency (SIT)
Note that the accumulated values are symmetric, i.e., $Accu_{ij}(\mathcal{G},H,t_{cur})=Accu_{ji}(\mathcal{G},H,t_{cur})$, reflecting the reciprocity of friendship.
@@
5.3. Comparative Enjoyment Degree (CED)
Users often engage in multiple groups in online social networks (Tan and Lee, 2015). They perceive greater cohesiveness in a specific group when it positively contributes to their enjoyment, as indicated by Items 7 and 9. This leads us to propose comparative enjoyment degree to measure how much more enjoyment a user gets from community interactions than those outside the community.
@@ 5.3. Comparative Enjoyment Degree (CED)
[Comparative Enjoyment Degree (CED)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and a user $u_{i}\in H$, the CED of $u_{i}$ at time $t_{cur}$ is defined as the difference between the enjoyment derived from interactions within and outside $H$ by time $t_{cur}$ :
@@
5.4. Group Interaction Preference (GIP)
For the GI-S scale, Item 11 captures the group’s preference for group interaction over individual posting, which can be formulated as group interaction preference (GIP). Let $\mathcal{E}_{H,inter}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},j\neq i,t\leq t_{cur}\}$ and $\mathcal{E}_{H}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},t\leq t_{cur}\}$ denote the set of interaction activities and all activities in $H$ up to $t_{cur}$, respectively, where $\mathcal{E}_{H,inter}^{t_{cur}}\subseteq\mathcal{E}_{H}^{t_{cur}}$. The GIP is defined as follows.
@@ 5.4. Group Interaction Preference (GIP)
$\mathbb{GIP}(\mathcal{G},H,t_{cur})=|\mathcal{E}_{H,inter}^{t_{cur}}|\big{/}|\mathcal{E}_{H}^{t_{cur}}|.$
Notice that GIP ranges from 0 to 1.
5.5. Group Interaction Density (GID)
Lastly, Item 13 emphasizes the interplay between member relationships and interaction frequency, which can be quantified using weighted density (Liu et al., 2009). In our context, we term this measure as group interaction density (GID).
@@ 5.5. Group Interaction Density (GID)
[Group Interaction Density (GID)] Given a graph $\mathcal{G}$, community $H$ and current time $t_{cur}$, GID measures the ratio of user interactions to the total possible interacting relationships within $H$ by $t_{cur}$:
$\mathbb{GID}(\mathcal{G},H,t_{cur})=|\mathcal{E}_{H,inter}^{t_{cur}}|\big{/}\left(|\mathcal{U}_{H}|\cdot(|\mathcal{U}_{H}|-1)\right).$
@@ 6.1. Dataset
Sentiment Analysis. The preprocessed tweets were then subjected to a sentiment analysis task to determine their sentiment scores. Following common practice, we treat it as a three-class classification problem (Chang et al., 2024). To ensure superior analysis results, we adopted the Llama3-8B, an open-source Large Language Model (LLM) from Meta (AI@Meta, 2024), and used a prompt-based approach. We adapted the prompt from (Kheiri and Karimi, 2023) to assign sentiment labels (see Figure 2). We mapped these labels to numerical values 1, 0, and -1, respectively. Both top_p and temperature were set to 0.01 for more deterministic responses. Only the tweets successfully assigned labels were retained.
@@ 6.2. Representative CS Algorithms
ALS identifies the $\beta$-temporal proximity core with the largest $\beta$ w.r.t. a query node in a temporal graph (Lin et al., 2024), integrating temporal proximity with structural cohesiveness seamlessly. It requires specifying the teleportation probability $\alpha$ when computing temporal proximity. WCF-CRC searches on dynamic networks that are represented as a sequence of time-variant weighted graph instances ($\mathbb{G}={G_{t_{1}},G_{t_{2}},\ldots,G_{t_{T}}}$) (Tang et al., 2022). It aims to find a Time Interval based ($\theta,k$)-core Reliable Community (CRC) with the highest reliability score, where a ($\theta,k$)-core in a $G_{t}$ is a $k$-core where each edge weight is at least $\theta$. CSD finds a $(k,l)$-core in directed graphs based on a query vertex $q$, where $k$ and $l$ denote the minimum in-degree and out-degree of subgraph nodes (Chen et al., 2021).
@@
Given a query vertex $q$ and a size constraint $[l,h]$, ST-Exa identifies a subgraph with the largest min-support while satisfying the size constraint. It is a parameter-free technique that eliminates the need to specify the cohesiveness constraint. Repeeling+ is a peeling-based algorithm for finding a maximal $D$-truss containing the query node in the snapshot digraph $G^{t}$ (Liao et al., 2024). $G^{t}$ is induced from a streaming directed graph using a sliding window length $\tau$ and a stride $\beta$. The algorithm’s parameters need to be tuned according to the dataset characteristics to ensure valid results. I2ACSM aims to search an attributed $pkd$-truss community that includes the query node in an Attributed and Weighted Graph (AWG) (Xie et al., 2021), ensuring that any vertex in the $k$-truss subgraph is within distance $d$ from the query node and the maximum path influence from the query node to any subgraph node exceeds $p$ (Xie et al., 2021).
@@ 6.4. Parameter Settings
We use an exponential decay function, commonly applied in fields such as biology, physics, and network analysis (Hobbie et al., 2007; Karagiannis et al., 2007; Stehlé et al., 2010), to calculate cohesiveness measures. The function is defined as $\Phi(t_{cur}-t)=e^{-\lambda(t_{cur}-t)}$, where $t_{cur}$ is the latest timestamp in each dataset. The decay rate $\lambda$ is selected from $\{0.0001,0.0005,0.001,0.005,0.01\}$, with $0.0001$ as the default. As users in the four datasets lack consistent information except their ids, we used only user_id as the node feature during CSGphormer training.
@@ 7. Experimental Evaluation
In this section, we conduct an in-depth evaluation of the representative CS algorithms. We report key results here. The reader may refer to (tech) for detailed results. We first assess the structural cohesiveness of their identified communities, followed by a detailed evaluation of their quality based on five psychology-informed cohesiveness measures. Learning-based algorithms were run on a Linux server with an Intel Xeon Platinum 8255C CPU (2.50GHz) and an RTX 3080 GPU (10GB memory). The remaining experiments were conducted on a Linux server with an Intel Xeon(R) Platinum 8352V CPU (2.10GHz) and 120GB memory. Recall that we do not focus on efficiency and scalability analysis here, as our goal is to understand the effectiveness of CS algorithms. The reader may refer to (Fang et al., 2020) for studies related to the former.
@@ 8. Reflections & Future Work
Despite these significant results, our evaluation has some limitations. Firstly, the design of our measures is constrained by sentiment analysis tools, and the simple three-class classification in our sentiment analysis may overlook nuanced sentiments inherent in specific comments. Secondly, our study is based on a single social network platform (i.e., X). However, given the fundamental limitations in the design of existing cohesiveness measures for human-interaction datasets (Section 4.2), we believe the key insights reported in this paper will also hold in other large online social networks. Thirdly, our study cannot use ground-truth communities for benchmarking because no public datasets are created based on cohesion theories.

### v2

@@ 1. Introduction
In summary, this work makes the following novel contributions. (a) It is the first study to assess the cohesiveness of social network communities generated by various community search methods, analyzed through the lens of cohesion theories in social psychology. (b) It introduces a set of psychology-informed cohesiveness measures for evaluation, which is unavailable in existing community search evaluation toolkits. (c) The experimental findings presented in this work are novel and have not been reported in previous studies.
@@ 3. The Community Search Problem
Problem 3.1.
[Community Search (CS) Problem] Given a graph $G=(V,E)$ and a set of query nodes $Q\subseteq V$, the community search problem seeks to find a community that contains $Q$ and satisfies the following properties:
@@
Traditional CS algorithms. Generally, traditional CS algorithms formulate a community by specifying its subgraph properties (Fang et al., 2020). The commonly adopted cohesiveness property is structural-based, including k-core (Sozio and Gionis, 2010; Cui et al., 2014; Barbieri et al., 2015; Yao and Chang, 2021), k-truss (Huang et al., 2014; Akbas and Zhao, 2017; Liu et al., 2020; Zhang et al., 2023), k-clique (Yuan et al., 2017), and k-ECC (Chang et al., 2015; Hu et al., 2016, 2017). These metrics are often referred to as structural cohesiveness metrics (structural metrics for brevity). Additionally, studies incorporate attribute information to formalize cohesiveness by refining existing structural metrics or proposing new attribute cohesiveness metrics alongside the structural ones. Examples of the former approach include attributed $pkd$-truss community (Xie et al., 2021), $(k,\theta)$-community (Miao et al., 2022), $(k,p)$-core (Lu et al., 2022), and $\beta$-temporal core (Lin et al., 2024). Examples of the latter include keyword cohesiveness (Fang et al., 2016), spatial cohesiveness (Fang et al., 2017; Chen et al., 2018), and query cohesiveness (Das et al., 2021).
@@ 4.1. Methodology
To measure this concept, a conceptual model has been proposed that identifies two key perceptions binding members to their group: a member’s perception of the group as a whole (labeled as group integration (GI)) and a member’s personal attraction to the group (labeled as individual attractions to group (ATG)). Each perception is further divided into task and social aspects, resulting in four correlated constructs that influence group cohesion: group integration-task (GI-T), group integration-social (GI-S), individual attractions to group-task (ATG-T), and individual attractions to group-social (ATG-S) (Carron et al., 1985). The Group Environment Questionnaire (GEQ) derived from this model consists of 18 items and has demonstrated its reliability and validity in various contexts (Carron et al., 1985; Blanchard et al., 2000; Anderson and Warner, 2017). The cohesion score for a group is calculated by averaging individual GEQ responses (Carron et al., 2003).
@@ 4.2. Comparison
Table 2 summarizes whether existing measures (introduced in Section 3) capture the cohesion features outlined in the adapted GEQ (Table 1). Observe that both ATG and GI items incorporate underlying structural information, yet in distinct ways. Moreover, ATG places greater emphasis on personal opinions, carrying more psychological insights that influence a user’s membership. Items are labeled “S” for structural information and “P” for psychological information. $\checkmark$”, $\triangle$”, and “$\times$” denote that an item is fully captured, partially captured, or not captured by existing measures, respectively.
@@ 5. Psychology-informed Cohesiveness Measures
5.1. Measures for ATG-S Items
Enjoyment Index (EI). Item 1 emphasizes the user’s enjoyment level during social interactions. Enjoyment is often measured concerning positive emotions or needs fulfillment during interactions (Syn and Oh, 2015; Kawamichi et al., 2016; Lin et al., 2009). Accordingly, we identify the enjoyment index as a cohesiveness indicator, quantified by the cumulative sentiments elicited during interactions.
@@ 5.1. Measures for ATG-S Items
Let $\mathcal{E}_{H,i\cdot}^{t_{cur}}=\{e_{xy}^{t}\mid e_{xy}^{t}\in\mathcal{E}_{H},\{x,y\}=\{i,j\},j\neq i,t\leq t_{cur}\}$ be activities where user $u_{i}$ interacts with other group members up to time $t_{cur}$. Assuming that more recent activities better capture user interests (Sarker et al., 2019), we weight them by a decay function $\Phi(\cdot)$ and define the enjoyment index as follows.
@@
Sentimental Interaction Tendency (SIT). Now consider Item 5 in Table 1. Having best friends in the group indicates cohesiveness, prompting the need to measure friendships among group members. Friendship generally develops through the exchange of information (Chan and Lo, 2014) and is characterized by reciprocity, closeness, and emotional attachment (Amichai-Hamburger et al., 2013; Paxton and Moody, 2003). Compared to ordinary friends, best friends share greater intimacy, reciprocity, and stability (Berndt, 1982; Roberto and Kimboko, 1989; Weisz and Wood, 2005). For simplicity, we propose sentimental interaction tendency to characterize friendships developed through cumulative information exchange, with higher values denoting closer ties.
@@
Note that $Accu_{ij}(\cdot)$ is symmetric, i.e., $Accu_{ij}(\mathcal{G},H,t_{cur})=\linebreak Accu_{ji}(\mathcal{G},H,t_{cur})$, reflecting the reciprocity of friendship.
@@
Comparative Enjoyment Degree (CED). Users often engage in multiple groups in online social networks (Tan and Lee, 2015). They perceive greater cohesiveness in a specific group when it positively contributes to their enjoyment, as indicated by Items 7 and 9. This leads us to propose comparative enjoyment degree to measure how much more enjoyment a user gets from community interactions than those outside the community.
@@
[Comparative Enjoyment Degree (CED)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and a user $u_{i}\in H$, the CED of $u_{i}$ at time $t_{cur}$ is defined as the difference between the enjoyment derived from interactions within and outside $H$ by $t_{cur}$ :
@@
5.2. Measures for GI-S Items
Group Interaction Preference (GIP). Item 11 captures the group’s preference for group interaction over individual posting, which can be formulated as group interaction preference (GIP). Let $\mathcal{E}_{H,inter}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},j\neq i,t\leq t_{cur}\}$ and $\mathcal{E}_{H}^{t_{cur}}=\{e_{ij}^{t}\mid e_{ij}^{t}\in\mathcal{E}_{H},t\leq t_{cur}\}$ denote the set of interaction activities and all activities in $H$ up to $t_{cur}$, respectively, where $\mathcal{E}_{H,inter}^{t_{cur}}\subseteq\mathcal{E}_{H}^{t_{cur}}$. The GIP ranges from 0 to 1 and is defined as follows.
@@ 5.2. Measures for GI-S Items
$\mathbb{GIP}(\mathcal{G},H,t_{cur})=\frac{|\mathcal{E}_{H,inter}^{t_{cur}}|}{|\mathcal{E}_{H}^{t_{cur}}|}.$
Group Interaction Density (GID). Lastly, Item 13 emphasizes the interplay between member relationships and interaction frequency. We quantify it by extending the weighted density (Liu et al., 2009) with a temporal normalization. Let the observation window be $W_{t_{cur}}=(t_{cur}-t_{0})/\tau_{t}$, where $t_{0}$ is the start of observation and $\tau_{t}$ is the time-unit (e.g., day, week, month). We call the resulting metric the group interaction density (GID).
@@
[Group Interaction Density (GID)] Given a graph $\mathcal{G}$, community $H$, current time $t_{cur}$, and observation window $W_{t_{cur}}$, GID measures the interaction frequency per possible user pair per unit time within $H$ by $t_{cur}$:
$\mathbb{GID}(\mathcal{G},H,t_{cur})=\frac{|\mathcal{E}_{H,inter}^{t_{cur}}|}{|\mathcal{U}_{H}|\cdot(|\mathcal{U}_{H}|-1)\cdot W_{t_{cur}}}.$
@@ 6.1. Dataset
Sentiment Analysis. The preprocessed tweets were then subjected to a sentiment analysis task to determine their sentiment scores. Following common practice, we treat it as a three-class classification problem (Chang et al., 2024). To ensure superior analysis results, we adopted the Llama3-8B, an open-source Large Language Model (LLM) from Meta (AI@Meta, 2024), and used a prompt-based approach. The prompt was adapted from (Kheiri and Karimi, 2023) to assign sentiment labels (see Figure 2), which were mapped to 1, 0, and -1, respectively. Both top_p and temperature were set to 0.01 for more deterministic responses. Only the tweets successfully assigned labels were retained.
@@ 6.2. Representative CS Algorithms
ALS identifies the $\beta$-temporal proximity core with the largest $\beta$ in a temporal graph (Lin et al., 2024), integrating temporal proximity with structural cohesiveness seamlessly. It requires specifying the teleportation probability $\alpha$ to compute temporal proximity. WCF-CRC searches on dynamic networks, represented as a sequence of time-variant weighted graph instances ($\mathbb{G}={G_{t_{1}},G_{t_{2}},\ldots,G_{t_{T}}}$) (Tang et al., 2022). It aims to find a Time Interval based ($\theta,k$)-core Reliable Community (CRC) with the highest reliability score, where a ($\theta,k$)-core in a $G_{t}$ is a $k$-core where all edge weights are at least $\theta$. CSD finds a $(k,l)$-core in the directed graph, where $k$ and $l$ denote the minimum in-degree and out-degree of subgraph nodes (Chen et al., 2021).
ST-Exa identifies a subgraph with the largest min-support based on the query node, while satisfying the size constraint $[l,h]$. It is parameter-free, eliminating the need to specify the cohesiveness constraint. Repeeling+ is a peeling-based algorithm for finding a maximal $D$-truss containing the query node in the snapshot digraph $G^{t}$ (Liao et al., 2024). $G^{t}$ is induced from a streaming directed graph using a sliding window length $\tau$ and a stride $\beta$. The algorithm’s parameters need to be tuned according to the dataset characteristics to ensure valid results. I2ACSM aims to search an attributed $pkd$-truss community that includes the query node in an Attributed and Weighted Graph (AWG) (Xie et al., 2021), ensuring that any vertex in the $k$-truss subgraph is within distance $d$ from the query node and the maximum path influence from the query node to any subgraph node exceeds $p$ (Xie et al., 2021).
@@ 6.4. Parameter Settings
We use an exponential decay function, commonly applied in fields such as biology, physics, and network analysis (Hobbie et al., 2007; Karagiannis et al., 2007; Stehlé et al., 2010), to calculate cohesiveness measures. The function is defined as $\Phi(t_{cur}-t)=e^{-\lambda(t_{cur}-t)}$, where $t_{cur}$ is the latest timestamp in each dataset. The decay rate $\lambda$ is selected from $\{0.0001,0.0005,0.001,0.005,0.01\}$, with $0.0001$ as the default. When calculating GID values, we note that the four datasets span different observation lengths and that no standard choice of $\tau_{t}$ exists. Since $W_{t_{cur}}$ uniformly scales all GID scores of communities within a dataset without affecting their relative ranking, we omit $W_{t_{cur}}$ for clarity and simplicity. As users in the four datasets lack consistent information except their ids, we used only user_id as the node feature during CSGphormer training.
Figure 3. The CHASE framework.
@@ 7. Experimental Evaluation
In this section, we conduct an in-depth evaluation of the representative CS algorithms. We first assess the structural cohesiveness of their identified communities, followed by a detailed evaluation of their quality based on five psychology-informed cohesiveness measures. Learning-based algorithms were run on a Linux server with an Intel Xeon Platinum 8255C CPU (2.50GHz) and an RTX 3080 GPU (10GB memory). The remaining experiments were conducted on a Linux server with an Intel Xeon(R) Platinum 8352V CPU (2.10GHz) and 120GB memory. Recall that we do not focus on efficiency and scalability analysis here, as our goal is to understand the effectiveness of CS algorithms. The reader may refer to (Fang et al., 2020) for studies related to the former.
@@ 8. Reflections & Future Work
Despite these significant results, our evaluation has some limitations. Firstly, the design of our measures is constrained by sentiment analysis tools, and the simple three-class classification in our sentiment analysis may overlook nuanced sentiments inherent in specific comments. Secondly, our study is based on a single social network platform (i.e., X). However, given the fundamental limitations in the design of existing cohesiveness measures for human-interaction datasets (Section 4.2), we believe the key insights reported in this paper will also hold in other large online social networks. Thirdly, our study cannot use ground-truth communities for benchmarking because no public ground-truth datasets are created based on cohesion theories.